# Riemannian Metric Learning: Closer to You than You Imagine

### v1

@@ Acknowledgments
for the tangent space. Grassmannians occur, for example, in the context of video based face recognition [93, 75, 78, 92]. Via the quotient construction, we obtain again a metric that descents from the Stiefel manifold and is simply the Euclidean metric for this description of the tangent space, i.e. $\mathfrak{g}_{P}(V,W)=\mathrm{tr}V^{T}W$. The corresponding Riemannian distance $\mathbf{d}_{\mathfrak{g}}$ is also called the Grassmanian distance. There is a range of further distances constructed on the Grassmannian, which are not-necessarily Riemannian distances. The most popular of these distances is the so-called projection distance $\mathbf{d}(P_{1},P_{2})\coloneqq\lVert P_{1}-P_{2}\rVert_{F}$ which is thus the Euclidean distance of matrices applied to the projections. It yields a computationally cheap local approximation of the Grassmanian distance, i.e. its differentiation yields the same metric as shown e.g. by Harandi et al. [76]. This projection distance has been used by a number of approaches to distance/metric learning on Grassmannians such as [75, 93]. Geometrically, the projection and the Grassmanian distance can also be described using the principal angles between the involved subspaces, which is also true for other proposed distances. For a list see, for example, [199, Table 2]. For a more general introduction into the geometry of the Grassmanian and the Stiefel manifold, see, for example, [21].
@@
The idea of the Grassmannian can be generalized by considering nonlinear, curved submanifolds of $\mathbb{R}^{n}$ instead of linear subspaces. To this end, one considers a base manifold $M$ with dimension smaller than $n$. Then one studies the space of all its immersions $\mathrm{Imm}(M,\mathbb{R}^{n})$ into $\mathbb{R}^{n}$ of sufficient regularity. It can be shown that $\mathrm{Imm}(M,\mathbb{R}^{n})$ is indeed a manifold under certain assumptions [87], albeit an infinite-dimensional one [^4] We do not discuss this here but the ideas from Section 2 can be generalized to infinite-dimensional spaces, see Appendix A.. When considering closed planar curves, i.e. $\mathrm{Imm}(S^{1},\mathbb{R}^{2})$, this manifold can be used, e.g., to model the boundaries of cells in microscopy images. Furthermore, surfaces immersed into $\mathbb{R}^{3}$ are used in computer graphics to model characters or similar and in medical imaging to model the shape of anatomical objects. The tangent space of $\mathrm{Imm}(M,\mathbb{R}^{n})$ consists of vector fields along the immersed surface on which various metrics on the space of immersions have been proposed. Popular choices include Sobolev metric of vector fields [15], metrics derived from elastic energies [83], and finally metrics that descend from metrics on the space of diffeomorphisms [3]. For the latter, one considers the action of diffeomorphisms on immersions by deforming the ambient space and introduces a metric on the Lie group of diffeomorphims of sufficient regularity that is compatible with this action.
@@
Nevertheless, considering data to lie on Riemannian manifolds allows us to encode additional structure of the data which is particularly interesting when working in data sparse regimes. In this case, this additional structure often helps to improve generalization. For example, van Veldhuizen [180] showed that using a curved, hyperspherical latent space within Variational Autoencoders improved performances compared to a flat Euclidean one. In the context of metric learning, we of course aim to imprint (some of) the structure of the data onto the metric. By providing a Riemannian metric on an Euclidean space, we also obtain a curved space and, by moving points to infinite distance, we can even encode some topology in the metric [41]. However, in some cases, having the manifold structure already as a strong prior can facilitate the learning process and improve performance, especially concerning generalization. These two aspects combined yield an intricate decision between how much structure to prescribe explicitly via the manifold and how much to learn via the metric which has to be made on a case-by-case basis and will be illuminated more in the following sections.
@@
Note that the integration relies on the metric volume $\text{vol}_{\mathfrak{g}}$ in such way that $\mathbf{G}_{\mathfrak{g}}$ is not a trivial extension of the affine-invariant metric. However, the smoothness of Riemannian metrics does not appear explicitly in $\mathbf{G}_{\mathfrak{g}}$. Even if working in the manifold $\mathcal{G}_{\mathcal{M}}$ might appear challenging, it turns out that a closed-formed expression is available for the geodesics which is given in [67, Theorem 3.2] even when the base manifold $\mathcal{M}$ is not compact. Nevertheless, the geodesics are expensive to compute. While this structure seems natural, it is not yet used to perform Riemannian metric learning as we will point out in Section 5.
@@
If a Riemannian metric learning method is used, at some point, the dataset $\mathsf{X}=(x_{i})_{i\in[N]}$ has to be generated according to an underlying metric $\mathfrak{g}^{*}$ related to a manifold $\mathcal{M}$, i.e. for any $i\in[N]$, $x_{i}$ is the realization of $X_{i}\sim Q_{\mathfrak{g}^{*}}^{i}$ where $(Q_{\mathfrak{g}^{*}}^{i})_{i\in[N]}$ are probability laws on the input space $\mathcal{X}$ and $\mathcal{M}\subset\mathcal{X}$. Note that $X_{i}$ is not necessarily in $\mathcal{M}$ because of possible additive noise $\varepsilon$ such that $X_{i}=M_{i}+\varepsilon$ where $M_{i}\in\mathcal{M}$. The dataset $\mathsf{X}$ can have a more complex structure, but we assume that $\mathsf{X}=(x_{i})_{i}$ here for the sake of simplicity and more complex datasets will be discussed in Section 4.2
Optimizing on the manifold $\mathcal{G}_{\mathcal{M}}$ of all possible infinitely differentiable Riemannian metric on $\mathcal{M}$ can be tractable in dimension 1 [121, Chapter 4-5], but most of times it is not. Thus, a family of parametrized Riemannian metrics $\mathcal{G}_{\theta}\subset\mathcal{G}_{\mathcal{M}}$ is chosen. The wealth of Riemannian metric learning as an optimization problem arise with the use of different mathematical objects, to name the most commons: the geodesics distance $\mathbf{d}_{\mathfrak{g}}$, the exponential map $\mathrm{Exp}_{\mathfrak{g}}$, the parallel transport $\mathfrak{T}_{\mathfrak{g}}$, the metric volume $\text{vol}_{\mathfrak{g}}$. Therefore, denoting by $\mathfrak{L}$ the objective function, the abstract form of Riemannian metric learning can be stated as follow :
@@
    The manifold of Riemannian metrics $\mathcal{G}_{\mathcal{M}}$ has a specific structure [67] (see (2.26)). They are not only vectors or constrained-matrices in $\mathcal{S}_{++}$, they are $\mathcal{S}_{++}$-valued smooth functions with a manifold input space $\mathcal{M}$. $\mathcal{G}_{\mathcal{M}}$ is both constrained and infinite dimensional. This makes it challenging to find suitable discretizations of it, which will be further discussed in Section 4.1.
@@
However, Riemannian metric learning not only offers inescapable problems, but also yield new opportunities for representation learning to leverage a large number of data in an unsupervised manner as we will see in Section 4. Before we how to learn a Riemannian, we will explain why by looking at the applications in the next section.
@@
First, we will present the applications related to the distance metric learning literature (Section 3.2.1), then the applications related to Riemannian metric learning (Section 3.2.2). When the learned distance $\mathbf{d}_{\mathfrak{g}}$ is induced by a non-linear Riemannian metric $\mathfrak{g}$, the application is presented in Section 3.2.2 instead of Section 3.2.1. We aim to showcase the richness of the Riemannian metric learning approach and to highlight the importance of each Riemannian object of interest, namely $(\mathbf{d}_{\mathfrak{g}},\mathrm{Exp}_{\mathfrak{g}},\mathfrak{T}_{\mathfrak{g}},\text{vol}_{\mathfrak{g}})$. Given that Riemannian metric learning is inherently technical and complex, it is crucial to ensure its relevance in the context of the application.
@@
Learning the metric of a metric space (see Section 2) can be seen as a fondational task which has its most prominent use in distance-based regression, classification (k-NN), clustering and dimension reduction. Instead of specifying a distance according to your own prior, you can learn relevant information from the dataset and introduce it in your distance.
@@
On Euclidean data, the paper [195] popularized metric learning for clustering and classification using nine dataset from the UC Irvine repository [^7] https://archive.ics.uci.edu (UCI) by extending k-means [1]. This classical repository is composed principally of tabular dataset having less than 3000 samples and 10 classes with low dimensional inputs ($d$<100). Then, it was followed by [161, 71, 70, 46, 192, 102, 185, 203, 37] outperforming each other chronologically regarding accuracy and time of computations on UCI datasets, Handwritten text recognition (MNIST [^8] https://yann.lecun.com/exdb/mnist/, USPS [^9] https://paperswithcode.com/dataset/usps) and Face recognition datasets such as Yales [17], PubFig [138], LFW [91].
@@
As discussed in Section 3.1, a Riemannian metric can be used for modeling through different objects: the distance $\mathbf{d}_{\mathfrak{g}}$, the exponential map $\mathrm{Exp}_{\mathfrak{g}}$ related to geodesics $\gamma$, the parallel transport $\mathfrak{T}_{\mathfrak{g}}$, the metric volume $\text{vol}_{\mathfrak{g}}$ or even sometimes the Ricci curvature of the manifold $R_{\mathfrak{g}}$ [114, 59]. In the following, we clarify and exemplify why these objects can be useful for representation learning, Hamiltonian learning, temporal longitudinal modeling, distribution trajectory inference, sampling and generative modeling, domain adaptation, and finally causal inference.
@@
The first paper we found with “Riemannian metric learning” in its title was by Peltonen et al. [143] in 2004 for exploratory analysis and dimension reduction. They aruged that, although Riemannian metric learning increased the computational burden, they used it since global distance would not have preserved the topology of the data space, which they deemed important for identifying subcategories of the classes within a dataset [143, p.3]. Other papers also used principles of Riemannian metric learning, without saying it, for instance Isomap [173, 2000] is arguably one of the most famous. In Isomap [173], the geodesic distance $\mathbf{d}_{\mathfrak{g}}$ is defined according to the k-nearest neighbors graph, which can cause robustness issues [12]. Recently, a variant of Isomap using parallel transport $\mathfrak{T}$ proposed by Budninskiy et al. [28] revealed to be more robust. Therefore, representations can be learned through Riemannian metric learning, not only by using distances but also by using connections on tangent bundles.
@@
Then, any standard method of dimension reduction (e.g. PCA) can be applied to these representations $(\log_{p}(x_{i}))_{i\in[N]}$ [80, 116, 181, 65]. This strategy can thus transform points $(x_{i})_{i}$ on a Riemannian manifold $\mathcal{M}$, where there is not necessarily a group structure, to vectors $(\log_{p}(x_{i}))_{i\in[N]}$ in a vector space $T_{p}M$, where there is a group structure. This opens the toolbox of Euclidean statistical methods, which is appealing to analyze complex objects such as shapes [179] as depicted in Figure 6 with mouse respiratory cycles.
Simpler than a normal coordinate chart is a global coordinate chart $\psi$ as used in Section 2 and implicitly in autoencoders [13] and variational autoencoders [105]. Here, the observations $y=\psi_{\theta}(z)$ are represented by low dimensional latent variables $z\in\mathcal{Z}$. This idea was further developed in [99, 63, 34, 8, 60, 48, 49] to enhance training and interpretability of the representation through visualization of geometric features, which might also encode task-related uncertainty [191, 171] as depicted in Figure 6. Once a latent space and its geometric structure $(\mathcal{Z},\mathfrak{g})$ are learned, different applications involving Riemannian distance, geodesics, or simply features representation can be performed. Geodesics in the latent space are especially appealing for temporal longitudinal modeling and distribution trajectory inference as detailed in the next sections.
Figure 6: On the left (A), an example taken from [65] (with permission) of normal coordinate chart representing shapes of respiratory cycles related to two mouse genotypes: colq and wt. On the right (B), an example taken from [171] (licensed under CC BY 4.0) of latent representation where the metric encodes uncertainty information related to an ensemble of decoders $\psi_{\theta}$.
@@
Before presenting the applications related to geodesics modeling, we include a discussion on the broader context of Hamiltonian learning. In physics, trajectories are often analyzed using Hamiltonians, which describe preserved energy along the trajectory. The Hamiltonian related to an object at position $q$ and momentum $p$ is denoted by $\mathcal{H}(q,p)$ and decomposes as $\mathcal{H}(q,p)=V(q)+p^{\top}\mathfrak{g}(q)^{-1}p/2$ with $V(q)$ being the potential energy at position $q$ and $p^{\top}\mathfrak{g}(q)^{-1}p/2$ the kinetic energy depending on the metric of the space. Recently, deep learning methods were suggested [176, 40] to model the underlying Hamiltonian as a neural network and infer it from observed trajectories. This is depicted on the left of Figure 7.
Neural Hamiltonians connect with Riemannian metric learning since geodesics are Hamiltonian trajectories for the special case of $V\equiv 0$ as mentioned in Section 2. When the kinectic energy is fixed within Neural Hamiltonians, i.e. $\mathfrak{g}$ constant, we only need to analyze the potential $V$ to get insight about the state space, which can enhance the interpretability of the Hamiltonian learning [167]. In the other way around, if $V$ is removed or fixed, learning the metric $\mathfrak{g}$ shed light not only on states $q$, but also on displacement $p$. It is not only about where your are, but where you want to go. In the next paragraph, we illustrate these ideas with Longitudinal modeling and Trajectory inference.
@@
In temporal longitudinal modeling [^10] This is not the general case of longitudinal modeling. In longitudinal modeling the outcome variable $Y(x)$ might depends on a covariate $x$ which is not a scalar as timepoints., the same individual $i$ is observed with features $(y_{i}^{j})$ at different time points $(t_{i}^{j})$ which can be irregularly sampled. The goal is to recover the underlying process $t\to Y_{i}(t)$ such that $Y_{i}(t_{i}^{j})=y_{i}^{j}+\varepsilon_{i}^{j}$ with $\varepsilon_{i}^{j}$ an i.i.d. noise. One approach is to model the individual trajectories $Y_{i}$ as geodesics, i.e. for any $t\in[0,1]$, $Y_{i}(t)=\mathrm{Exp}_{q_{i}}(tv_{i})$ with $(q_{i},v_{i})\in~\mathcal{M}\times T\mathcal{M}$. The idea is that geodesics generalized straight lines which make them both flexible and simple enough to avoid overfitting on noisy datasets. Geodesics learning was also used to model trajectories on rotations manifold to interpolate animations [109], human or robot motion captures [16, 175], image registration [163], and taxi routes in traffic jams [23, 148]. We should note that geodesic learning can also occur in exotic application such as a two player game where the goal of the first player is to install a surveillance system and the goal of the second is to reach a target area without being detected [134].
Figure 7: On the left, an example of Hamiltonian learning taken from [40] (licensed under CC BY 4.0). Neural Hamiltonian learning is compared to a naive neural networks encoding the system by its differential equation. Note that predicted baseline trajecories are thicker because of the larger uncertainty of estimation on the derivatives compared to the Hamiltonian. On the right, an example taken from [72] (with permission) where the geodesics are learned to mimic AD patient trajectories. HIPP is the cerebral volume of the Hippocampus and MMSE refers to the Mini Mental Status Examination score. The scores were reversed and normalized in order to be increasing in $[0,1]$. The learned exponential shape in figure b is consistent with clinical observations, but is not present when considering a hand-cafted logistic metric [159] showing the interest of learning the Riemannian metric from scratch. The tubular green and blue curves are the exp-paralellization of the average geodesic in red for different values of space shifts $w$ (3.3), it represents the inter-indiviudal variability of trajectories along the mean trajectory.
@@
for any $t\in\mathsf{I}$. This shows how parallel transport and the exponential can be combined when optimizing a Riemannian metric. Note that the exp-parallelization of a geodesic is not necessarily a geodesic. An illustration of the transformation is given on the right of Figure 7. Alzheimer’s disease does not start at the same age for everyone and does not evolve at the same speed. This generates inter-patient temporal variability. Thus, it is crucial to compose an exp-parallelization with an individual time warp $\psi_{i}\colon\mathsf{I}\mapsto\mathbb{R}$ as $Y_{i}=\operatorname*{ExPar}(w_{i})(\mathrm{Exp}_{q_{0}}(v_{0}\cdot),\psi_{i})$ with $(q_{0},v_{0},w_{i})\in~\mathcal{M}\times T_{q_{0}}\mathcal{M}\times T_{q_{0}}\mathcal{M}$. Parameters related to time ($\psi_{i}$) and space ($q_{0},v_{0},w_{i}$) are statisticaly identifiable as long as $||v_{0}||=1$ since the geodesics are parametrized by arc length (Section 2). This property is particularly interesting to disentangle time and space variability in the analysis.

### v2

@@ Acknowledgments
for the tangent space. Grassmannians occur, for example, in the context of video based face recognition [96, 78, 81, 95]. Via the quotient construction, we obtain again a metric that descents from the Stiefel manifold and is simply the Euclidean metric for this description of the tangent space, i.e. ${\mathfrak{g}}_{P}(V,W)=\mathrm{tr}V^{T}W$. The corresponding Riemannian distance $\mathbf{d}_{\mathfrak{g}}$ is also called the Grassmanian distance. There is a range of further distances constructed on the Grassmannian, which are not-necessarily Riemannian distances. The most popular of these distances is the so-called projection distance $\mathbf{d}(P_{1},P_{2})\coloneqq\lVert P_{1}-P_{2}\rVert_{F}$ which is thus the Euclidean distance of matrices applied to the projections. It yields a computationally cheap local approximation of the Grassmanian distance, i.e. its differentiation yields the same metric as shown e.g. by Harandi et al. [79]. This projection distance has been used by a number of approaches to distance/metric learning on Grassmannians such as [78, 96]. Geometrically, the projection and the Grassmanian distance can also be described using the principal angles between the involved subspaces, which is also true for other proposed distances. For a list see, for example, [206, Table 2]. For a more general introduction into the geometry of the Grassmanian and the Stiefel manifold, see, for example, [23].
@@
The idea of the Grassmannian can be generalized by considering nonlinear, curved submanifolds of $\mathbb{R}^{n}$ instead of linear subspaces. To this end, one considers a base manifold $M$ with dimension smaller than $n$. Then one studies the space of all its immersions $\mathrm{Imm}(M,\mathbb{R}^{n})$ into $\mathbb{R}^{n}$ of sufficient regularity. It can be shown that $\mathrm{Imm}(M,\mathbb{R}^{n})$ is indeed a manifold under certain assumptions [90], albeit an infinite-dimensional one [^4] We do not discuss this here but the ideas from Section 2 can be generalized to infinite-dimensional spaces, see Appendix A.. When considering closed planar curves, i.e. $\mathrm{Imm}(S^{1},\mathbb{R}^{2})$, this manifold can be used, e.g., to model the boundaries of cells in microscopy images. Furthermore, surfaces immersed into $\mathbb{R}^{3}$ are used in computer graphics to model characters or similar and in medical imaging to model the shape of anatomical objects. The tangent space of $\mathrm{Imm}(M,\mathbb{R}^{n})$ consists of vector fields along the immersed surface on which various metrics on the space of immersions have been proposed. Popular choices include Sobolev metric of vector fields [16, 17, 141], metrics derived from elastic energies [86], and finally metrics that descend from metrics on the space of diffeomorphisms [4]. For the latter, one considers the action of diffeomorphisms on immersions by deforming the ambient space and introduces a metric on the Lie group of diffeomorphims of sufficient regularity that is compatible with this action. Elasticity-derived or Sobolev metrics include parameters weighing different contributions, e.g. different orders of derivatives or different gradient directions (e.g. normal vs. tangential). These parameters are classically derived from physical material properties, however, in the context of Riemannian metric learning, they yield a parametrized family of metrics to optimize over [17].
@@
Nevertheless, considering data to lie on Riemannian manifolds allows us to encode additional structure of the data which is particularly interesting when working in data sparse regimes. In this case, this additional structure often helps to improve generalization. For example, van Veldhuizen [187] showed that using a curved, hyperspherical latent space within Variational Autoencoders improved performances compared to a flat Euclidean one. Theoretically, Suzuki et al. [177, 176] proved that embeddings in hyperbolic space offer better statistical generalization guarantees than those in Euclidean space for hierarchical tree-like data.
In the context of metric learning, we of course aim to imprint (some of) the structure of the data onto the metric, not only on the manifold. By providing a Riemannian metric on an Euclidean space, we also obtain a curved space and, by moving points to infinite distance, we can even encode some topology in the metric [44]. However, in some cases, having the manifold structure already as a strong prior can facilitate the learning process and improve performance, especially concerning generalization. These two aspects combined yield an intricate decision between how much structure to prescribe explicitly via the manifold and how much to learn via the metric which has to be made on a case-by-case basis and will be illuminated more in the following sections.
@@
Note that the integration relies on the metric volume $\text{vol}_{\mathfrak{g}}$ in such way that $\mathbf{G}_{\mathfrak{g}}$ is not a trivial extension of the affine-invariant metric. However, the smoothness of Riemannian metrics does not appear explicitly in $\mathbf{G}_{\mathfrak{g}}$. Even if working in the manifold $\mathcal{G}_{\mathcal{M}}$ might appear challenging, it turns out that a closed-formed expression is available for the geodesics which is given in [70, Theorem 3.2] even when the base manifold $\mathcal{M}$ is not compact. Nevertheless, the geodesics are expensive to compute. While this structure seems natural, it is not yet used to perform Riemannian metric learning as we will point out in Section˜5.
@@
If a Riemannian metric learning method is used, at some point, the dataset $\mathsf{X}=(x_{i})_{i\in[N]}$ has to be generated according to an underlying metric $\mathfrak{g}$$^{*}$ related to a manifold $\mathcal{M}$, i.e. for any $i\in[N]$, $x_{i}$ is the realization of $X_{i}\sim Q_{{\mathfrak{g}}^{*}}^{i}$ where $(Q_{{\mathfrak{g}}^{*}}^{i})_{i\in[N]}$ are probability laws on the input space $\mathcal{X}$ and $\mathcal{M}\subset\mathcal{X}$. Note that $X_{i}$ is not necessarily in $\mathcal{M}$ because of possible additive noise $\varepsilon$ such that $X_{i}=M_{i}+\varepsilon$ where $M_{i}\in\mathcal{M}$. The dataset $\mathsf{X}$ can have a more complex structure, but we assume that $\mathsf{X}=(x_{i})_{i}$ here for the sake of simplicity and more complex datasets will be discussed in Section˜4.2
Optimizing on the manifold $\mathcal{G}_{\mathcal{M}}$ of all possible infinitely differentiable Riemannian metric on $\mathcal{M}$ can be tractable in dimension 1 [124, Chapter 4-5], but most of times it is not. Thus, a family of parametrized Riemannian metrics $\mathcal{G}_{\theta}\subset\mathcal{G}_{\mathcal{M}}$ is chosen. The wealth of Riemannian metric learning as an optimization problem arise with the use of different mathematical objects, to name the most commons: the geodesics distance $\mathbf{d}_{\mathfrak{g}}$, the exponential map $\mathrm{Exp}_{\mathfrak{g}}$, the parallel transport $\mathfrak{T}_{\mathfrak{g}}$, the metric volume $\text{vol}_{\mathfrak{g}}$. Therefore, denoting by $\mathfrak{L}$ the objective function, the abstract form of Riemannian metric learning can be stated as follow :
@@
    The manifold of Riemannian metrics $\mathcal{G}_{\mathcal{M}}$ has a specific structure [70] (see (2.26)). They are not only vectors or constrained-matrices in $\mathcal{S}_{++}$, they are $\mathcal{S}_{++}$-valued smooth functions with a manifold input space $\mathcal{M}$. $\mathcal{G}_{\mathcal{M}}$ is both constrained and infinite dimensional. This makes it challenging to find suitable discretizations of it, which will be further discussed in Section˜4.1.
@@
However, Riemannian metric learning not only offers inescapable problems, but also yield new opportunities for representation learning to leverage a large number of data in an unsupervised manner as we will see in Section˜4. Before we how to learn a Riemannian, we will explain why by looking at the applications in the next section.
@@
First, we will present the applications related to the distance metric learning literature (Section˜3.2.1), then the applications related to Riemannian metric learning (Section˜3.2.2). When the learned distance $\mathbf{d}_{\mathfrak{g}}$ is induced by a non-linear Riemannian metric $\mathfrak{g}$}, the application is presented in Section˜3.2.2 instead of Section˜3.2.1. We aim to showcase the richness of the Riemannian metric learning approach and to highlight the importance of each Riemannian object of interest, namely $(\mathbf{d}_{\mathfrak{g}},\mathrm{Exp}_{\mathfrak{g}},\mathfrak{T}_{\mathfrak{g}},\text{vol}_{\mathfrak{g}})$. Given that Riemannian metric learning is inherently technical and complex, it is crucial to ensure its relevance in the context of the application.
@@
Learning the metric of a metric space (see Section˜2) can be seen as a fondational task which has its most prominent use in distance-based regression, classification (k-NN), clustering and dimension reduction. Instead of specifying a distance according to your own prior, you can learn relevant information from the dataset and introduce it in your distance.
@@
On Euclidean data, the paper [202] popularized metric learning for clustering and classification using nine dataset from the UC Irvine repository [^7] https://archive.ics.uci.edu (UCI) by extending k-means [2]. This classical repository is composed principally of tabular dataset having less than 3000 samples and 10 classes with low dimensional inputs (dd<100). Then, it was followed by [166, 74, 73, 49, 199, 105, 192, 210, 40] outperforming each other chronologically regarding accuracy and time of computations on UCI datasets, Handwritten text recognition (MNIST [^8] https://yann.lecun.com/exdb/mnist/, USPS [^9] https://paperswithcode.com/dataset/usps) and Face recognition datasets such as Yales [19], PubFig [142], LFW [94].
@@
As discussed in Section˜3.1, a Riemannian metric can be used for modeling through different objects: the distance $\mathbf{d}_{\mathfrak{g}}$, the exponential map $\mathrm{Exp}_{\mathfrak{g}}$ related to geodesics $\gamma$, the parallel transport $\mathfrak{T}_{\mathfrak{g}}$, the metric volume $\text{vol}_{\mathfrak{g}}$ or even sometimes the Ricci curvature of the manifold $R_{\mathfrak{g}}$ [117, 62]. In the following, we clarify and exemplify why these objects can be useful for representation learning, Hamiltonian learning, temporal longitudinal modeling, distribution trajectory inference, sampling and generative modeling, domain adaptation, and finally causal inference.
@@
The first paper we found with “Riemannian metric learning” in its title was by Peltonen et al. [147] in 2004 for exploratory analysis and dimension reduction. They aruged that, although Riemannian metric learning increased the computational burden, they used it since global distance would not have preserved the topology of the data space, which they deemed important for identifying subcategories of the classes within a dataset [147, p.3]. Other papers also used principles of Riemannian metric learning, without saying it, for instance Isomap [180, 2000] is arguably one of the most famous. In Isomap [180], the geodesic distance $\mathbf{d}_{\mathfrak{g}}$ is defined according to the k-nearest neighbors graph, which can cause robustness issues [13]. Recently, a variant of Isomap using parallel transport $\mathfrak{T}$ proposed by Budninskiy et al. [31] revealed to be more robust. Therefore, representations can be learned through Riemannian metric learning, not only by using distances but also by using connections on tangent bundles.
@@
Then, any standard method of dimension reduction (e.g. PCA) can be applied to these representations $(\log_{p}(x_{i}))_{i\in[N]}$ [83, 119, 188, 68]. This strategy can thus transform points $(x_{i})_{i}$ on a Riemannian manifold $\mathcal{M}$, where there is not necessarily a group structure, to vectors $(\log_{p}(x_{i}))_{i\in[N]}$ in a vector space $T_{p}M$, where there is a group structure. This opens the toolbox of Euclidean statistical methods, which is appealing to analyze complex objects such as shapes [186] as depicted in Figure˜6 with mouse respiratory cycles.
Simpler than a normal coordinate chart is a global coordinate chart $\psi$ as used in Section˜2 and implicitly in autoencoders [14] and variational autoencoders [108]. Here, the observations $y=\psi_{\theta}(z)$ are represented by low dimensional latent variables $z\in\mathcal{Z}$. This idea was further developed in [102, 66, 37, 9, 63, 51, 52] to enhance training and interpretability of the representation through visualization of geometric features, which might also encode task-related uncertainty [198, 178] as depicted in Figure˜6. Once a latent space and its geometric structure $(\mathcal{Z},{\mathfrak{g}})$ are learned, different applications involving Riemannian distance, geodesics, or simply features representation can be performed. Estimates of manifold curvature are even employed in the study of brain neural activation signals [1]. Geodesics in the latent space are especially appealing for temporal longitudinal modeling and distribution trajectory inference as detailed in the next sections.
Figure 6: On the left (A), an example taken from [68] (with permission) of normal coordinate chart representing shapes of respiratory cycles related to two mouse genotypes: colq and wt. On the right (B), an example taken from [178] (licensed under $C$ BY 4.0) of latent representation where the metric encodes uncertainty information related to an ensemble of decoders $\psi_{\theta}$.
@@
Before presenting the applications related to geodesics modeling, we include a discussion on the broader context of Hamiltonian learning. In physics, trajectories are often analyzed using Hamiltonians, which describe preserved energy along the trajectory. The Hamiltonian related to an object at position $q$ and momentum $p$ is denoted by $\mathcal{H}(q,p)$ and decomposes as $\mathcal{H}(q,p)=V(q)+p^{\top}{\mathfrak{g}}(q)^{-1}p/2$ with $V(q)$ being the potential energy at position $q$ and $p^{\top}{\mathfrak{g}}(q)^{-1}p/2$ the kinetic energy depending on the metric of the space. Recently, deep learning methods were suggested [183, 43] to model the underlying Hamiltonian as a neural network and infer it from observed trajectories. This is depicted on the left of Figure˜7.
Neural Hamiltonians connect with Riemannian metric learning since geodesics are Hamiltonian trajectories for the special case of $V\equiv 0$ as mentioned in Section˜2. When the kinectic energy is fixed within Neural Hamiltonians, i.e. $\mathfrak{g}$} constant, we only need to analyze the potential $V$ to get insight about the state space, which can enhance the interpretability of the Hamiltonian learning [172]. In the other way around, if $V$ is removed or fixed, learning the metric $\mathfrak{g}$} shed light not only on states $q$, but also on displacement $p$. It is not only about where your are, but where you want to go. In the next paragraph, we illustrate these ideas with Longitudinal modeling and Trajectory inference.
@@
In temporal longitudinal modeling [^10] This is not the general case of longitudinal modeling. In longitudinal modeling the outcome variable $Y(x)$ might depends on a covariate $x$ which is not a scalar as timepoints., the same individual $i$ is observed with features $(y_{i}^{j})$ at different time points $(t_{i}^{j})$ which can be irregularly sampled. The goal is to recover the underlying process $t\to Y_{i}(t)$ such that $Y_{i}(t_{i}^{j})=y_{i}^{j}+\varepsilon_{i}^{j}$ with $\varepsilon_{i}^{j}$ an i.i.d. noise. One approach is to model the individual trajectories $Y_{i}$ as geodesics, i.e. for any $t\in[0,1]$, $Y_{i}(t)=\mathrm{Exp}_{q_{i}}(tv_{i})$ with $(q_{i},v_{i})\in \mathcal{M}\times T\mathcal{M}$. The idea is that geodesics generalized straight lines which make them both flexible and simple enough to avoid overfitting on noisy datasets. Geodesics learning was also used to model trajectories on rotations manifold to interpolate animations [112], human or robot motion captures [18, 182], image registration [168, 141], and taxi routes in traffic jams [25, 152]. We should note that geodesic learning can also occur in exotic application such as a two player game where the goal of the first player is to install a surveillance system and the goal of the second is to reach a target area without being detected [137].
Figure 7: On the left, an example of Hamiltonian learning taken from [43] (licensed under $C$ BY 4.0). Neural Hamiltonian learning is compared to a naive neural networks encoding the system by its differential equation. Note that predicted baseline trajecories are thicker because of the larger uncertainty of estimation on the derivatives compared to the Hamiltonian. On the right, an example taken from [75] (with permission) where the geodesics are learned to mimic AD patient trajectories. HIPP is the cerebral volume of the Hippocampus and MMSE refers to the Mini Mental Status Examination score. The scores were reversed and normalized in order to be increasing in $[0,1]$. The learned exponential shape in figure b is consistent with clinical observations, but is not present when considering a hand-cafted logistic metric [164] showing the interest of learning the Riemannian metric from scratch. The tubular green and blue curves are the exp-paralellization of the average geodesic in red for different values of space shifts $w$ (3.3), it represents the inter-indiviudal variability of trajectories along the mean trajectory.
@@
for any $t\in\mathsf{I}$. This shows how parallel transport and the exponential can be combined when optimizing a Riemannian metric. Note that the exp-parallelization of a geodesic is not necessarily a geodesic. An illustration of the transformation is given on the right of Figure˜7. Alzheimer’s disease does not start at the same age for everyone and does not evolve at the same speed. This generates inter-patient temporal variability. Thus, it is crucial to compose an exp-parallelization with an individual time warp $\psi_{i}\colon\mathsf{I}\mapsto\mathbb{R}$ as $Y_{i}=\operatorname*{ExPar}(w_{i})(\mathrm{Exp}_{q_{0}}(v_{0}\cdot),\psi_{i})$ with $(q_{0},v_{0},w_{i})\in \mathcal{M}\times T_{q_{0}}\mathcal{M}\times T_{q_{0}}\mathcal{M}$. Parameters related to time ($\psi_{i}$) and space ($q_{0},v_{0},w_{i}$) are statisticaly identifiable as long as $v_{0}||=1$ since the geodesics are parametrized by arc length (Section˜2). This property is particularly interesting to disentangle time and space variability in the analysis.